OpenStudy (geometry_hater):

Perimeter of the rhombus MNPQ is 48 and the perimeter of the parallelogram NPRS, which has a common side with the rhombus, is 60. Which is the difference between two distinct sides of the parallelogram? 10 9 8 6

OpenStudy (anonymous):

6

OpenStudy (anonymous):

18-12=6

OpenStudy (anonymous):

the side of the rhombus is 48/4=12 then one couple sides of the parm is 12 so the other couple is 36/2=18 18-12=6